neglecting friction and air resistance what force id required to accelerate an automobile weighing 5.0x10n from 60km/hr to 85km/hr in 5minutes


1
Expert's answer
2021-06-29T01:52:04-0400
"F=ma\\F=\frac{W}{g}\frac{v_f-v_i}{t}\\F=\frac{5000}{9.8}\frac{85-60}{5\cdot 60\cdot3.6}=12\ N"